I am trying to put the points (-7,2) and (-1,7) in general form.
Answer by checkley71(8403)

FIRST YOU NEED THE SLOPE (Y2-Y1)/(X2-X1)
(7-2)/(-1+7)
5/6 IS THE SLOPE.
NOW SUBSTITUTE A SET OF X,Y POINTS AND SOLVE FOR b (Y INTERCEPT)
7=5/6*-1+b
7=-5/6+b
b=7 5/6
THUS THE EQUATION OF THIS LINE FORMULA IS
Y=5/6X+7 5/6